OpenStudy (anonymous):

-√80 how do i do this?

OpenStudy (anonymous):

I assume you just need to factor it out so the best way to do it is figure out a multiple that's a perfect square and then it's just a multiple of that and the square root of the other number, such as squrt(125) is 5 * squrt(5) since 125 = 5 * 25.

OpenStudy (anonymous):

well the instructions say that i need to simplify and that all vaiables need to represent nonnegative numbers...

OpenStudy (anonymous):

so how do i take the negative away?

OpenStudy (anonymous):

Take out the multiple of sqrt(1) which is the imaginary unit, i as a multiple of your simplified answer, my friend.

OpenStudy (anonymous):

im so confused

OpenStudy (phi):

If this is \[ -\sqrt{80} \] then the best you can do is factor out a perfect square 80 = 16*5 or 4^2 * 5 so you get \[-\sqrt{16*5} =-4\sqrt{5} \]